OpenStudy (anonymous):

partial derivatives of:

OpenStudy (anonymous):

\[f(x,y) = \int\limits_{y}^{x}\cos(t ^{2})dt\]

OpenStudy (zarkon):

use the fundamental theorem of calculus

OpenStudy (anonymous):

what's that?

OpenStudy (anonymous):

i still don't understand how to do it

OpenStudy (zarkon):

\[\frac{\partial}{\partial x}\int\limits_{y}^{x}\cos(t ^{2})dt=\cos(x^2)\] \[\frac{\partial}{\partial y}\int\limits_{y}^{x}\cos(t ^{2})dt=-\frac{\partial}{\partial y}\int\limits_{x}^{y}\cos(t ^{2})dt=-\cos(y^2)\]

OpenStudy (zarkon):

let \[F(t)=\int f(t)dt\] then \[\int\limits_{y}^{x}f(t)dt=F(x)-F(y)\] thus \[\frac{\partial}{\partial x}\int\limits_{y}^{x}f(t)dt=\frac{\partial}{\partial x}(F(x)-F(y))=F'(x)=f(x)\] and \[\frac{\partial}{\partial y}\int\limits_{y}^{x}f(t)dt=\frac{\partial}{\partial y}(F(x)-F(y))=-F'(y)=-f(y)\]

OpenStudy (anonymous):

so wrt x it's just cos(x^2)?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

and wrt y it's -cos(x^2)?

OpenStudy (zarkon):

\[-\cos(y^2)\]

OpenStudy (anonymous):

ah got cha! thanks!